OpenStudy (anonymous):

Help!! I'm doing homework right now and I'm really struggling with it. could anyone walk me though f(x)=3|2x-1|+x find f(-4)

OpenStudy (anonymous):

Try simplifying this. \[3(2x + 1)\] Can be simplified to \[5x + 5\] by multiplying the inside of the brackets by 3.

OpenStudy (anonymous):

okay that makes since can you keep explaining? I'm kinda getting it

OpenStudy (anonymous):

This makes the sum \[f(x) = (5x + 5) + x\] You can simplify it further by removing the '+x' by putting it inside the brackets. That gives you \[F(x) = 6x + 5\] Now try removing x from the first side. That'll give you f. Remember, what ever you do to one side you must do to the other.

OpenStudy (anonymous):

so to get f by itself would I multiply x or 5?

OpenStudy (anonymous):

I mean subtract

OpenStudy (anonymous):

F(x) is F times X, so you remove the x by dividing both sides by x.

OpenStudy (anonymous):

okay so if I do that I get f=6x+4 ? I'm sorry this is so confusing.

OpenStudy (anonymous):

Huh? Ok. 6x divided by x is 6. does that help?

OpenStudy (anonymous):

OH so f=6+5?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

NO WAIT GOD IT'S A FUNCTION

OpenStudy (anonymous):

ERASE EVERYTHING I JUST SAID

OpenStudy (anonymous):

GOD I'm SO SORRY

OpenStudy (anonymous):

it's okay haha

OpenStudy (anonymous):

Right. This is much easier, actually. Don't actually erase everything, because the simplification still stands You just need to work out \[6x + 5\] if X is -4.

OpenStudy (anonymous):

that was my bad I should of told you😂

OpenStudy (anonymous):

so would I do 6•-4+5?

OpenStudy (anonymous):

No, I was a muppet and completely forgot F(x) means x is input thus function.

OpenStudy (anonymous):

Yeah, exactly. For real and proper this time :)

OpenStudy (anonymous):

okay so is that all since I found -4's function or is there more?(:

OpenStudy (anonymous):

Yup, that's all. For functions, if you need to find f and you have the input, you just substitute the input for x.

OpenStudy (anonymous):

thank you so much this makes total since now! you are a life saver

OpenStudy (anonymous):

On the plus side, you now know how to solve both\[fx = 3(2x+1) + x\] and \[f(x) 3(2x + 1) +x\] which are for some reason completely different.

OpenStudy (anonymous):

No problem! Thank you for forgiving my muppetosity :)

OpenStudy (anonymous):

so my answer to this problem is just 6•-4+5

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

okay thank you again

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

GOD NOT AGAIN I'M SO STUPID It doesn't simplify to 5x + 5. It's 6x + 3. So \[F(-4) = (6x + 3) + x\] equals \[F(-4) = (7 \times -4) +3\] So F(-4) = -25. This time I'm pretty sure I'm right, but give me a minute and I'm sure I'll find out I'm counting in the wrong language or something.